OpenStudy (anonymous):

simplify the square root of -4 + the square root of -12

OpenStudy (anonymous):

\[\sqrt{-4} + \sqrt{-12}\]

OpenStudy (anonymous):

\[2i + 3.46410162i = \]5.46410162 i

OpenStudy (anonymous):

It'd be an imaginary number

OpenStudy (anonymous):

thank you!